OpenStudy (anonymous):

what is y intercept of 12x+4y=10

OpenStudy (tester97):

all you have to do is re arange this equation :) seems easy right?

OpenStudy (tester97):

Isolate y on one side of the equal side and everything else on the other.

OpenStudy (anonymous):

it doesnt seem easy

OpenStudy (tester97):

Ok let me try to show you in steps

OpenStudy (anonymous):

kk

OpenStudy (tester97):

\[12x+4y=10\] All of this is divisible by two so lets simplify that\[6x+2y=5\] Now we move the 2y over.\[2y+5=6x\]Next we move the 5 to the other side \[2y=6x-5\]Next step is do divide both sides by 2 so it would be \[y=3x-\frac{ 5 }{ 2 }\]And viola thats your answer :)

OpenStudy (anonymous):

thanks(:

OpenStudy (tester97):

Welcome :)